# Retain or Reframe? A Computational Framework for the Analysis of Framing in News Articles and Reader Comments

### v1

@@ Abstract
When a news article describes immigration as an "economic burden" or a "humanitarian crisis," it selectively emphasizes certain aspects of the issue. Although framing shapes how the public interprets such issues, audiences do not absorb frames passively but actively reorganize the presented information. While this relationship between source content and audience response is well-documented in the social sciences, NLP approaches often ignore it, detecting frames in articles and responses in isolation. We present the first computational framework for large-scale analysis of framing across source content (news articles) and audience responses (reader comments). Methodologically, we refine frame labels and develop a framework that reconstructs dominant frames in articles and comments from sentence-level predictions, and aligns articles with topically relevant comments. Applying our framework across eleven topics and two news outlets, we find that frame reuse in comments correlates highly across outlets, while topic-specific patterns vary. We release a frame classifier that performs well on both articles and comments, a dataset of article and comment sentences manually labeled for frames, and a large-scale dataset of articles and comments with predicted frame labels. [^1] We released our at: https://anonymous.4open.science/r/FrAC-A8E3/; and fine-tuned frame classifier: https://huggingface.co/mattdr/sentence-frame-classifier
@@ 1 Introduction
Framing research has long recognized that how news content is presented fundamentally shapes public understanding of issues (Entman, 1993; Chong and Druckman, 2007). However, audiences do not passively absorb frames but rather "actively filter, sort and reorganize information in personally meaningful ways in the process of constructing an understanding of public issues" (Neuman et al., 1992, pg.77). Such a constructionist understanding emphasizes that frames cannot be fully understood without considering the interaction between source and receiver (Scheufele, 1999; Gorp, 2007; Hubner and Dixon, 2023). This is particularly evident on news websites with comment threads that create immediate opportunities for readers to publicly respond to framed content — a paradigm shift from one-way communication to participatory discourse where audiences actively engage with and reconstruct news content (Risch and Krestel, 2020).
@@
To address this gap, we propose a computational framework that investigates at scale the relationship between how information is framed in news articles and how readers respond to such framed content in associated comments. Drawing from the constructionist understanding of framing (Gorp, 2007) and using an established framing taxonomy (Boydstun et al., 2014; Card et al., 2015), we examine both source content and audience responses as interconnected elements of a broader communicative process (Hubner and Dixon, 2023).
Throughout this paper, we use media framing effects to describe the relation between news article framing and that of reader comments, as an umbrella term that covers two phenomena: frame retention occurs when readers adopt the dominant frame used in the source article in their comments, while reframing occurs when readers shift to a different dominant frame in their comments. Figure 1 illustrates both, with an example.
Our large-scale analysis across eleven diverse topics and two news outlets [^2] The New York Times (USA) and The Globe and Mail (Canada) reveals systematic patterns of media framing effects. We find that (a) frame retention rates vary by source document frame, and this variation correlates highly across the two outlets, with value-laden frames being more often reframed; (b) retention patterns vary systematically by topic, and (c) when reframing, readers predominantly shift discussions toward political interpretation regardless of the original article frame.
@@ 2 Related Work
Cognitive research established that news frames significantly influence audience interpretation using controlled experimental designs, where participants were assigned to different framing conditions of the same issue and asked to report their thoughts. Valkenburg et al. (1999) examined how news stories framed as conflict, human interest, attribution of responsibility, or economic consequences affected readers, finding that participants generally adopted news frames. However, Price et al. (1997) showed that frames not only guide interpretation but also stimulate new thoughts and feelings not explicitly present in the original content. Nelson and Oxley (1999) and Shen (2004) showed that framing primarily influences the importance of existing beliefs rather than changing the beliefs themselves, with frames most effective when resonating with pre-existing schemas. While producing valuable insights, these studies were restricted to small-scale experimental settings and narrow demographic representation. This limited their generalizability to framing effects in natural media environments with direct audience responses.
@@
Other studies have found no evidence of retention. Holton et al. (2014) analyzed online health coverage across major outlets and found no evidence of frame retention in comment threads, noting a disconnect between thematic news frames and predominantly episodic comment frames. Muñiz et al. (2015) examined coverage of an oil company expropriation in Spanish online outlets and found that media frames did not produce corresponding effects in reader comments. Similarly, Zhou and Moy (2007) found no framing effects in news about Chinese sociopolitical incidents. Finally, in climate change coverage, Koteyko et al. (2013) found minimal retention of frames in reader comments on online articles from UK outlets, while Hubner and Dixon (2023) showed that retweets sharing climate news rarely replicated the original frames.
@@
NLP approaches to framing have evolved from exploratory topic-modeling (DiMaggio et al., 2013) to structured classification tasks (for a review, see Ali and Hassan (2022)). The Media Frame Corpus (MFC, Card et al. (2015)), based on Boydstun et al. (2014), has been particularly influential, spurring various frame classification methods including probabilistic approaches (e.g. Johnson and Goldwasser (2018)), neural networks (e.g. Naderi and Hirst (2017)), and pre-trained language models (e.g. Khanehzar et al. (2021)). These computational approaches have analyzed frames at different granularities: headline level (Liu et al., 2019a; Akyürek et al., 2020), document level (Field et al., 2018; Ji and Smith, 2017; Piskorski et al., 2023), span level (Khanehzar et al., 2021), sentence level (Naderi and Hirst, 2017; Hartmann et al., 2019), and paragraph level (Roy and Goldwasser, 2020). Some studies have investigated the generalizability of the MFC to other domains and languages: Khanehzar et al. (2019) found reduced performance when applying fine-tuned BERT-based classifiers to different English contexts (Australian parliamentary speeches), while Daffara et al. (2025) tested cross-linguistic transfer using both fine-tuned pre-trained language models and zero-shot LLM classification on Portuguese news articles. However, these studies either focused on cross-linguistic transfer, or on the diagnosis of the problem of cross-topic and genre generalization but fell short of offering a solution to the problem, which we do here.
@@ 3 Research Questions
The gaps in existing research motivate three fundamental questions about media framing effects in readers comments:
  1. 1.
    To what extent are frames retained vs changed? Previous studies have shown mixed results. We investigate at scale whether article frames consistently influence comment frames, and whether this influence varies systematically across different types of frames.
  2. 2.
    How do framing effects vary across topics? Topic-specific studies cannot reveal whether framing effects represent universal phenomena or topic-specific patterns. We examine whether framing effects are consistent or vary based on the topic being discussed.
  3. 3.
    When readers do not retain the dominant frame of the article, how do they reframe it? Understanding reframing patterns is crucial for comprehending how audiences actively reconstruct media messages. We examine systematic tendencies in how readers shift framing when they reject source frames.
@@ 4.1 Data
Our framework requires a data set of source documents $d_{i}$ (e.g. news articles, headlines, tweets, etc.) each paired with a set of response documents $R_{d_{i}}=\{r_{1},r_{2},\ldots\}$ (e.g. comments, retweets, blog posts, etc.) directly linked to the source ($D=\{(d_{i},R_{d_{i}})\}_{i=1}^{N}$). For the purposes of this study, we use two existing data sets of news articles with their associated comments:
@@
We note that the framework generalizes to any corpus where source documents are mapped to audience responses.
@@ 4.2 Framework Structure
The framework consists of three core components that work in concert to analyze framing dynamics between articles and comments (Figure 1):
@@
    Dominant Frame Detection: In line with prior work in cognitive and communication studies (Section 2), we identify and compare a single, dominant frame (defined as the most recurring one; Section 6.2) in source documents and responses. We thus require a method to determine the dominant frame in each input document.
@@
    Topic Identification: We assign each source document and reader response to a topic, for two reasons. First, assigning topics to articles allows us to perform analysis of framing effects over multiple topics, advancing over previous, topic-specific studies (Section 2) and potentially explaining their divergent results. Second, we want to capture framing effects only for responses where commentators stay on topic with the source, rather than digress into a different topic (with potentially a different frame, which can lead to underestimation of retention rates).
Below, we first introduce a new corpus that we annotated for the frame classification subtask (Section 5). Section 6 explains and evaluates our implementations of the three framework components introduced above. Finally, we apply the framework to construct a data set (Section 6.4) that we use to analyze media framing effects (Section 7).
@@ 5.1 Frame Labels and Granularity
We devise a robust frame label set, building on the Media Frames Corpus (MFC). The MFC is a foundational data set in computational framing analysis based on Boydstun et al. (2014) theoretical framework operationalizing framing through 15 topic-agnostic frame categories. The original MFC corpus Card et al. (2015) consists only of news articles; however, more recently the framework was successfully applied to other types of documents, including online forum comments Hartmann et al. (2019), suggesting suitability for our task that involves both long, formal documents and short, less formal responses.
@@
However, it has been noted (e.g., Ali and Hassan (2022); Daffara et al. (2025)) that some MFC frames are poorly defined and difficult to distinguish, leading to low inter-annotator agreement and model performance. First, to apply the MFC labels across our two text types (articles and comments), we need to define a granularity that is meaningful across both of them. While most of NLP studies on MFC looked at article-level dominant frames in the MFC, it is not feasible for comments (or other response documents), which are often brief and composed of single sentences. We therefore chose to use the sentence level – the smallest shared textual unit between the two types of text. [^5] Section 6.2 explains how we derive a dominant document frame from sentence-level frames.
@@
Second, to improve robustness, we isolated highly confused frame labels in the original MFC data set. To do so, we first identified spans-level MFC annotations (of ${\geq}3$ words) to which two or more annotators assigned different frame labels. We examined such inconsistent annotations using confusion matrices – both for all MFC data and separately for each topic included in the MFC corpus. We took a conservative approach by only merging frame labels that were confused consistently across all topics, [^6] For example, “Health and Safety” and “Legality, constitutionality and jurisprudence” labels are frequently assigned to the same spans in “Tobacco” and “Gun Control”, but not in other topics. In this case, we conclude that the confusion is due to the topic specifics rather than meaning of the label, and keep such semantically distinct labels separate. resulting in two merges of two frames (Table 1). To further improve robustness of the label set, we removed the two least frequent classes (“External regulation and reputation” and “Capacity and resources”) which accounted for less than 0.7% of the data. We finally obtained 9 frame categories, plus a generic “Other” frame to account for instances that do not match any of our nine frames. Table 1 lists the frames, with full definitions in Appendix Table 8.
@@
1) Legality and Crime, 2) Political and Policies+, 3) Economic, 4) Health and Safety, 5) Cultural Identity, 6) Public Opinion, 7) Morality, 8) Fairness and Equality, 9) Security and Defense, 10) Other.
@@
Table 1: The FrAC frame label set. Merged MFC labels "Legality, constitutionality and jurisprudence" and "Crime and punishment". + Merged MFC labels "Policy prescription and evaluation" and "Political".
@@ 5.2 Human Annotation
To validate our frame label set (and model predictions) we collect human annotation on a subset of the SOCC corpus (Section 4.1).
@@
We recruited six native English speakers as annotators, all with a background in NLP or political science and journalism. [^8] This study was approved by Human Ethics Committee (Referance No. 2025-32561-67793-4), and has been taken out to the according ethical standards. Annotators were compensated with vouchers over 50 USD, well above local minimum wage requirements. Each annotator was presented with 100 sentences from the SOCC corpus, drawn evenly from articles and comments, 20% of the sentences overlapped with another annotator to assess agreement. One additional annotator (an author of this paper) annotated all sentences (so all sentences in the resulting data set had at least two annotations). We adopted a multi-label annotation setup. Annotators were presented one input sentence at a time, and were asked to select at least one out of our 9 frames. If no frame matched, the category “Other” should be selected.
@@
As annotators could assign multiple labels to each sentence, we assessed inter-annotator agreement in two ways. Following prior work Arora et al. (2025), we use label-wise Krippendorff’s $\alpha$ to assess per-label agreement, and the Jaccard index to measure the averaged overlap in label sets between annotators. The averaged Krippendorff’s $\alpha$ across all labels and annotators indicates moderate to high agreement ($\alpha=0.76$), with some variance across frame labels (min for Morality with $\alpha=0.63$, max for Economic with $\alpha=0.89$). However even the category with lowest $\alpha$ reaches moderate agreement. The Jaccard index of 0.79 confirms this trend, suggesting a substantial overlap across label sets. This shows that a more robust set of labels applied at a sentence level allows to achieve substantial agreement despite the purported subjectivity of the task.
@@
To construct a gold standard for model evaluation, we applied majority voting among annotators for each sentence. A clear consensus (where either two or three annotators chose the same label) was achieved for 532 cases (where in 500 cases a single label was assigned, while in 32 all annotators have chosen two consistent labels). Sentences where annotators disagreed (45) were adjudicated by two authors of the paper, who independently re-assessed the cases of contention, choosing the label out of 9 categories. We kept only the cases where both adjudicators agreed with one of the original annotations, discarding 13 sentences. This process resulted in a validation data set of 556 sentences, which we use in Section 6.1 to assess model performance.

### v2

@@ Abstract
When a news article describes immigration as an "economic burden" or a "humanitarian crisis," it selectively emphasizes certain aspects of the issue. Although this framing shapes how the public interprets such issues, audiences do not absorb frames passively but actively reorganize the presented information. While this relationship between source content and audience response is well-documented in the social sciences, NLP approaches often ignore it, detecting frames in articles and responses in isolation. We present the first computational framework for large-scale analysis of framing across source content (news articles) and audience responses (reader comments). Methodologically, we refine frame labels and develop a framework that reconstructs dominant frames in articles and comments from sentence-level predictions, and aligns articles with topically relevant comments. Applying our framework across eleven topics and two news outlets, we find that frame reuse in comments correlates highly across outlets, and that readers often selectively engage with frames of the articles. We release a frame classifier that performs well on both articles and comments, a dataset of article and comment sentences manually labeled for frames, and a large-scale dataset of articles and comments with predicted frame labels. [^1] We released our data at: https://github.com/mattguida/FrAC; and fine-tuned frame classifier: https://huggingface.co/mattdr/sentence-frame-classifier
@@ 1 Introduction
Framing research has long recognized that how news content is presented fundamentally shapes public understanding of issues (Entman, 1993; Chong and Druckman, 2007). However, audiences do not passively absorb frames but rather "actively filter, sort and reorganize information in personally meaningful ways in the process of constructing an understanding of public issues" (Neuman et al., 1992, pg.77). The constructionist understanding of framing emphasizes that frames cannot be fully understood without considering the interaction between source and receiver (Scheufele, 1999; Gorp, 2007; Hubner and Dixon, 2023), as part of an interactive and dynamic process in which frames activate and reasonate (or fail to) with personal schemas through active meaning construction and intepretation. This is particularly evident on news websites with comment threads that create immediate opportunities for readers to publicly respond to framed content – a paradigm shift from one-way communication to participatory discourse where audiences actively engage with and reconstruct news content (Risch and Krestel, 2020).
@@
If framing is truly a process in which frames are actively constructed and interpreted by an audience (Scheufele, 1999; Gorp, 2007; Hubner and Dixon, 2023), then the audience’s response is not merely a posited outcome, but a fundamental part of the analysis that has been largely overlooked or observed at a limited scale. To address this gap, we propose a computational framework that investigates at scale the relationship between how information is framed in news articles and how readers respond to such framed content in associated comments. Using an established framing taxonomy (Boydstun et al., 2014; Card et al., 2015), we examine both source content and audience responses as interconnected elements of a broader communicative process (Hubner and Dixon, 2023).
Throughout this paper, we use the phrase media framing effects to describe the relation between news article framing and that of reader comments, as an umbrella term that covers two phenomena. Frame retention occurs when readers adopt the dominant frame [^2] We define the dominant frame as the one that covers the most sentences of an article, but at least 40% (see Section 6.2). used in the source article in their comments. Reframing occurs when readers shift to a different frame in their comments: either a minor frame in the article (selective reframing), or a completely new, unmentioned frame (complete reframing). Figure 1 illustrates this with an example.
Our large-scale analysis across eleven diverse topics and two news outlets (The New York Times and The Globe and Mail) reveals systematic patterns of media framing effects. We find that (a) frame retention rates vary by source document frame, and this variation correlates highly across the two outlets, with value-laden frames being more often reframed; (b) retention patterns vary systematically by topic, and (c) audiences engage in selective reframing, adopting secondary frames rather than completely accepting or rejecting primary frames.
@@ 2 Related Work
Cognitive research has established that news frames significantly influence audience interpretation using controlled experimental designs, where participants were assigned to different framing conditions of the same issue and asked to report their thoughts. Valkenburg et al. (1999) examined how news stories framed as conflict, human interest, attribution of responsibility, or economic consequences affected readers, finding that participants generally adopted news frames. However, Price et al. (1997) showed that frames not only guide interpretation but also stimulate new thoughts and feelings not explicitly present in the original content. Nelson and Oxley (1999) and Shen (2004) showed that framing primarily influences the importance of existing beliefs rather than changing them, with frames being the most effective when they resonate with pre-existing schemas. These studies were restricted to small-scale experimental settings and narrow demographic representation, and thus could not generalize to natural media environments with direct audience responses.
@@
Other studies have found no evidence of retention. Holton et al. (2014) analyzed online health coverage across major outlets and found no evidence of frame retention in comment threads. Muñiz et al. (2015) examined coverage of an oil company expropriation in Spanish online outlets and found that media frames did not produce corresponding effects in reader comments. Similarly, Zhou and Moy (2007) found no framing effects in news about Chinese sociopolitical incidents. Finally, in climate change coverage, Koteyko et al. (2013) found minimal retention of frames in reader comments on online articles from UK outlets, while Hubner and Dixon (2023) showed that retweets sharing climate news rarely replicated the original frames.
@@
NLP approaches to framing have evolved from exploratory topic-modeling (DiMaggio et al., 2013) to classification tasks (for a review, see Ali and Hassan (2022)). The Media Frame Corpus (MFC, Card et al. (2015)), based on Boydstun et al. (2014), has been particularly influential, spurring various frame classification methods including probabilistic approaches (e.g.,Johnson and Goldwasser (2018)), neural networks (e.g., Naderi and Hirst (2017)), and pre-trained language models (e.g., Khanehzar et al. (2021)). These computational approaches have analyzed frames at different granularities: headline level (Liu et al., 2019a; Akyürek et al., 2020), document level (Field et al., 2018; Ji and Smith, 2017; Piskorski et al., 2023), span level (Khanehzar et al., 2021), sentence level (Naderi and Hirst, 2017; Hartmann et al., 2019), and paragraph level (Roy and Goldwasser, 2020). Some studies have investigated the generalizability of the MFC to other domains and languages: Khanehzar et al. (2019) found reduced performance when applying fine-tuned BERT-based classifiers to different English contexts (Australian parliamentary speeches), while Daffara et al. (2025) tested cross-linguistic transfer using both fine-tuned pre-trained language models and zero-shot LLM classification on Portuguese news articles. However, these studies either focused on cross-linguistic transfer or on the diagnosis of the problem of cross-topic and genre generalization. They fell short of offering a solution to the problem, which we do here.
@@ 3 Research Questions
Our goal is to characterize and quantify the constructivist view on framing at scale. We formulate research questions that have been raised in the social sciences, albeit on a small scale often regarding a single topic and a specific framing inventory which makes it difficult to generalize results across scenarios. With orders of magnitude more data that covers a range of topics and is analyzed with a unified framing inventory, we are well positioned to answer the following research questions.
RQ1: To what extent are dominant article frames retained in reader comments?
We first focus on the extent to which the most prominent frame in the article is reflected in the reader comments. From a constructivist point of view, this assesses the extent to which a chosen article frame sticks with the readership.
Next, we address the fact that prior work was often topic-specific and led to inconclusive results about the extent of frame retention. Here, we can quantify the effect of topics on frame retention within a unified framework. We ask
RQ2: How does frame retention vary across topics?
Finally, we examine how readers how audiences reconstruct media messages when they do reframe (i.e., they choose not to retain the dominant article frame). Our framework allows us to paint a more nuanced picture on reframing than previous work. First, in line with prior studies we quantify how often readers adopt a frame that was never used in the article (complete reframing). In addition, we also explore how often readers shift from the dominant article frame to a secondary article frame (selective reframing). This leads to our final research question:
RQ3: If readers reject the dominant article frame, how do they reframe?
We next introduce our framework and methodology, before we address our research question in Section 7.
@@ 4.1 Data
Our framework requires a data set of source documents $d_{i}$ (e.g., news articles, headlines, tweets, etc.) each paired with a set of response documents $R_{d_{i}}=\{r_{1},r_{2},\ldots\}$ (e.g., comments, retweets, blog posts, etc.) directly linked to the source ($D=\{(d_{i},R_{d_{i}})\}_{i=1}^{N}$). For the purpose of this study, we use two existing data sets of news articles with their associated comments:
@@
This pairing of source documents with response documents directly operationalizes the constructionist understanding of framing. In our analysis, news articles represent the ’source framing’ – how journalists and media outlets choose to present issues through particular frames. Reader comments represent the ’audience response’ – how readers actively interpret, accept, modify, or reject those frames based on their own cognitive schemas and values. We focus exclusively on top-level comments that directly respond to articles, excluding nested replies within comment threads, as we want to capture direct audience responses to the original framed content.
@@ 4.2 Framework Structure
The framework consists of three core components that work in concert to analyze framing dynamics between articles and comments (Figure˜1):
@@
    Dominant Frame Detection: In line with prior work in cognitive and communication studies (Section 2), we identify and compare a single, dominant frame (defined as the most recurring one) in source documents and responses. We thus require a method to determine the dominant frame in each input document. We define the dominant frame as the most frequently occurring sentence-level frame within a document (see Section 6.2 for details). In doing so, we can distinguish the retention of the primary article frame from selective reframing into a secondary frame of the original article.
@@
    Topic Identification: We assign each source document and reader response to a topic, for two reasons. First, assigning topics to articles allows us to perform analysis of framing effects over multiple topics, advancing over previous, topic-specific studies (Section˜2) and potentially explaining their divergent results. Second, we want to capture framing effects only for responses where commentators stay on topic with the source, rather than digress into a different topic (with potentially a different frame, which can lead to underestimation of retention rates).
Below, we first introduce a new corpus that we annotated for the frame classification subtask (Section˜5). Section˜6 explains and evaluates our implementations of the three framework components introduced above. Finally, we apply the framework to construct a data set (Section˜6.4) that we use to analyze media framing effects (Section˜7).
@@ 5.1 Frame Labels and Granularity
We devise a robust frame label set, building on the Media Frames Corpus (MFC), a foundational data set in computational framing analysis based on the Boydstun et al. (2014)’s theoretical framework operationalizing framing through 15 generic frame categories. The original MFC corpus Card et al. (2015) consists only of news articles; however, more recently the framework has been successfully applied to other types of documents, including online forum comments Hartmann et al. (2019), suggesting suitability for our task that involves both long, formal documents and short, less formal responses.
@@
First, to apply the MFC labels across articles and comments, we need to define a granularity that is meaningful for both of them. While most NLP studies on the MFC modeled article-level dominant frames, this is not feasible for reader responses which are often brief and composed of single sentences. We therefore chose the sentence level – the smallest shared textual unit between the two types of text – as our basic unit of analysis. [^5] Section 6.2 explains how we derive a dominant document frame from sentence-level frames.
Second, we improve the robustness of the MFC label set as it has been found (e.g., Ali and Hassan (2022); Daffara et al. (2025)) that some MFC frames are poorly defined and difficult to distinguish, leading to low inter-annotator agreement and model performance.
@@
To do so, we isolated highly confused frame labels in the original MFC data set by first identifying span-level MFC annotations (of ${\geq}3$ words) to which two or more annotators assigned different frame labels. We examined such inconsistent annotations using confusion matrices – both for all MFC data and separately for each topic included in the MFC corpus. We took a conservative approach by only merging frame labels that were confused consistently across all topics, [^6] For example, “Health and Safety” and “Legality, constitutionality and jurisprudence” labels are frequently assigned to the same spans in “Tobacco” and “Gun Control”, but not in other topics. In this case, we conclude that the confusion is due to the topic specifics rather than meaning of the label, and keep such semantically distinct labels separate. resulting in two merges of two frames (Table˜1). To further improve robustness of the label set, we removed the two least frequent classes (“External regulation and reputation” and “Capacity and resources”) which accounted for less than 0.7% of the data. We finally obtained 9 frame categories, plus a generic “Other” frame to account for instances that do not match any of our nine frames. This category captures both genuinely unframed content and framed content that falls outside our nine established categories. Table 1 lists the frames, with the definitions following those of 10, as depicted in the Appendix by Table 10.
1) Legality and Crime, 2) Political and Policies+,
@@
3) Economic, 4) Health and Safety, 5) Cultural Identity, 6) Public Opinion, 7) Morality, 8) Fairness and Equality, 9) Security and Defense, 10) Other.
@@
Table 1: The FrAC frame label set. Label mergers: $\{$"Legality, constitutionality and jurisprudence", "Crime and punishment"}\}; ${=}\{$"Policy prescription and evaluation", "Political"}\}.
@@ 5.2 Human Annotation
To validate our frame label set (and model predictions) we collected human annotation on a subset of the SOCC corpus (Section˜4.1).
@@
We recruited six native English speakers as annotators, all with a background in NLP or political science and journalism. [^8] This study was approved by Human Ethics Committee (Referance No. 2025-32561-67793-4) and has been taken out to the according ethical standards. Annotators were compensated with vouchers over 50 USD, well above local minimum wage requirements. Each annotator was presented with 100 sentences from the SOCC corpus, drawn evenly from articles and comments, 20% of the sentences overlapped with another annotator to assess agreement. One additional annotator (an author of this paper) annotated all sentences (so all sentences in the resulting data set had at least two annotations). We adopted a multi-label annotation setup. Annotators were presented one input sentence at a time, and were asked to select at least one out of our 9 frames. If no frame matched, the category “Other” should be selected.
@@
Given our multi-label setup, we assessed inter-annotator agreement in two ways. Following prior work Arora et al. (2025), we use label-wise Krippendorff’s $\alpha$ to assess per-label agreement, and the Jaccard index to measure the averaged overlap in label sets between annotators. The averaged Krippendorff’s $\alpha$ across all labels and annotators indicates moderate to high agreement ($\alpha=0.76$), with some variance across frame labels (Table 2), however even the frame category with the lowest $\alpha$ reaches moderate agreement. We also compute an averaged Jaccard index of 0.79 which confirms this trend, suggesting a substantial overlap across label sets. This shows that a more robust set of labels applied at a sentence level allows to achieve substantial agreement despite the purported subjectivity of the task.
Label $\alpha$ Label $\alpha$
Economic 0.89 Political / Policy 0.78
Health / Safety 0.87 Cultural Identity 0.76
Legality / Crime 0.82 Other 0.66
Security / Defense 0.79 Fairness / Equality 0.65
Public Opinion 0.79 Morality 0.63
Table 2: Krippendorff’s $\alpha$ per label, averaged across annotation sessions and sorted in descending order.
@@
To construct a gold standard for model evaluation, we applied majority voting among annotators for each sentence. A clear consensus (where either two or three annotators chose the same label) was achieved for 532 cases (where in 500 cases a single label was assigned, while in 32 all annotators have chosen two consistent labels). Sentences where annotators disagreed (45) were adjudicated by two authors of the paper, who independently re-assessed the cases of contention, choosing the label out of 9 categories. We kept only the cases where both adjudicators agreed with one of the original annotations, discarding 13 sentences. This process resulted in a validation data set of 556 sentences, which we use in Section˜6.1 to assess model performance.